OpenStudy (beccaboo333):

Write the equation of the line that passes through (2, 4) and (1, -3) in slope-intercept form.

OpenStudy (beccaboo333):

A. y = 7x - 10 B. y = 7x + 10 C. y = 7x + 18 D. y = 7x - 18

ganeshie8 (ganeshie8):

first find the slope m, use slope formula

ganeshie8 (ganeshie8):

(2, 4) and (1, -3) x1 y1 x2 y2

ganeshie8 (ganeshie8):

slope, m = \(\large \frac{y2-y1}{x2-x1}\)

ganeshie8 (ganeshie8):

plug the numbers and simplify, wat do u get

OpenStudy (beccaboo333):

and this is why im going to fail math... i cant understand it

OpenStudy (beccaboo333):

im sorry

ganeshie8 (ganeshie8):

Answer : A. y = 7x - 10

ganeshie8 (ganeshie8):

we just need to try a bit + be patient i guess :)

OpenStudy (beccaboo333):

thanks :)

OpenStudy (beccaboo333):

3

OpenStudy (beccaboo333):

other questions